OpenStudy (anonymous):

Find the slope of the line passing through the points (-2, 9) and (5, 3). A. -6/7 B. 4 C. -7/6 D. 1/4

sam (.sam.):

\[\huge Slope=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

sam (.sam.):

\[\frac{3-9}{5--2}=-\frac{6}{7}\]